# In search of the electron-phonon contribution to total energy

### v1

@@ Acknowledgements.
The authors acknowledge useful discussions with Dr. Matteo Giantomassi and Prof. Matthieu Verstraete. S. P. acknowledges support from the Fonds de la Recherche Scientifique de Belgique (FRS-FNRS). This work was supported by the Fonds de la Recherche Scientifique - FNRS under Grants number T.0183.23 (PDR) and T.W011.23 (PDR-WEAVE). It is an outcome of the Shapeable 2D magnetoelectronics by design project (SHAPEme, EOSProjectNo. 560400077525) that has received funding from the FWO and FRS-FNRS under the Belgian Excellence of Science (EOS) program. This publication was supported by the Walloon Region in the strategic axe FRFS-WEL-T. Computational resources have been provided by the PRACE award granting access to MareNostrum4 at Barcelona Supercomputing Center (BSC), Spain and Discoverer in SofiaTech, Bulgaria (OptoSpin project id. 2020225411), by the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the FRS-FNRS under Grant No. 2.5020.11 and by the Walloon Region, as well as computational resources awarded on the Belgian share of the EuroHPC LUMI supercomputer.

### v2

@@ Acknowledgements.
The authors acknowledge useful discussions with Dr. Matteo Giantomassi, Prof. Matthieu Verstraete, and participants to the online ALPS meetings. S. P. acknowledges support from the Fonds de la Recherche Scientifique de Belgique (FRS-FNRS). This work was supported by the Fonds de la Recherche Scientifique - FNRS under Grants number T.0183.23 (PDR) and T.W011.23 (PDR-WEAVE). It is an outcome of the Shapeable 2D magnetoelectronics by design project (SHAPEme, EOSProjectNo. 560400077525) that has received funding from the FWO and FRS-FNRS under the Belgian Excellence of Science (EOS) program. This publication was supported by the Walloon Region in the strategic axe FRFS-WEL-T. Computational resources have been provided by the PRACE award granting access to MareNostrum4 at Barcelona Supercomputing Center (BSC), Spain and Discoverer in SofiaTech, Bulgaria (OptoSpin project id. 2020225411), by the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the FRS-FNRS under Grant No. 2.5020.11, by the Tier-1 supercomputer of the Walloon Region (Lucia) with infrastructure funded by the Walloon Region under the grant agreement n°1910247, and by the Belgian share of the EuroHPC LUMI supercomputer.